Given P(E) = 0.25 P(F)=0.6 and P(E u F ) = 0.7 find P(E n F)
1
Expert's answer
2020-10-13T19:03:07-0400

"P(E\cup F)=P(E)+P(F)-P(E\cap F)"

"P(E\cap F)=P(E)+P(F)-P(E\cup F)=0.25+0.6-0.7=0.15."